Calculate the total force exerted on the outside of a circular submarine window of diameter 30-cm at a depth of 1200 meters.

1
Expert's answer
2021-12-16T11:34:53-0500
"F=pA=\rho ghA=\rho gh\frac{\pi d^2}{4}=831265\text{ N}."